OpenStudy (anonymous):

please help <3 last test ever :) only 10 question so please help !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks <3

OpenStudy (bahrom7893):

is this really the last one? xD let me take a look

OpenStudy (bahrom7893):

A = initial amount = 3000

OpenStudy (anonymous):

yes i believe so .

OpenStudy (bahrom7893):

P = Ae^(kt) P = 3000e^(kt)

OpenStudy (bahrom7893):

P(6) = 3,600 3600=3000e^(k*6)

OpenStudy (bahrom7893):

36 = 30e^(6k) 6 = 5e^(6k)

OpenStudy (bahrom7893):

take Ln on both sides: Ln6 = ln[5e^(6k)] use Ln properties to simplify

OpenStudy (bahrom7893):

Ln6 = Ln5 + lne^(6k) ln6 = ln5 + 6k(lne) ln6 - ln5 = 6k since lne =1

OpenStudy (bahrom7893):

6k = 0.1823

OpenStudy (bahrom7893):

k = 0.1823/6 = 0.03

OpenStudy (bahrom7893):

Now put that back into P: P = 3000e^(kt) P = 3000e^(0.03t) Now what is this equal to after 13 hours?

OpenStudy (bahrom7893):

P(13) = 3000e^(0.03*13) = 4430.94 to the nearest whole that's like 4431

OpenStudy (amistre64):

you got to what: ln(P/A)/t = k right?

OpenStudy (bahrom7893):

I can't break this down any more than this..

OpenStudy (anonymous):

sorry i was in the bathroom and yes i understand :)

OpenStudy (anonymous):

yea omg i get it :) it was correct :D

OpenStudy (anonymous):

so all you do is replacing :) what u kno.

OpenStudy (bahrom7893):

yea

OpenStudy (anonymous):

thanks :) u wanna help me with the other 10 or no :( ?

OpenStudy (bahrom7893):

work this out posting the steps here, what do u do first?

OpenStudy (bahrom7893):

use initial pop = 197mln

OpenStudy (anonymous):

ok give me a sec :) im gonna try to solve it :)